# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ D.4.3 Comparaison of upper bounds in nonconvex settings
Let $n,f\in\mathbb{N}^{*},f<\frac{n}{2}$ and $g_{A},g_{B},g_{C}$ three collinear vectors in $\mathbb{R}^{d},d\geq 1$. Assume we observe a pool of $n$ vectors, consisting of $\alpha$ copies of $g_{A}$, $\beta$ copies of $g_{B}$ and $\gamma$ copies of $g_{C}$, where $\alpha,\beta,\gamma\in\{0,\ldots,n\}$, $\alpha+\beta+\gamma=n$. We abstract each subset $S\subseteq\{1,\ldots,n\}$, composed of $|S|=n-f$ vectors from our pool of vectors as $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that $a+b+c=n-f$ and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. With this notation, we have:
@@
Let $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that $a+b+c=n-f$ and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. Then $\overline{g}_{S}=\frac{1}{n-f}(ag_{A}+bg_{B}+cg_{C})$, $g_{A}-\overline{g}_{S}=\frac{b}{n-f}(g_{A}-g_{B})+\frac{c}{n-f}(g_{A}-g_{C})$, $g_{B}-\overline{g}_{S}=\frac{a}{n-f}(g_{B}-g_{A})+\frac{c}{n-f}(g_{B}-g_{C})$ and $g_{C}-\overline{g}_{S}=\frac{a}{n-f}(g_{C}-g_{A})+\frac{b}{n-f}(g_{C}-g_{B})$. We already note that the outer product of vectors $g_{A}$, $g_{B}$ and $g_{C}$ commute because they are collinear. We denote $(*)={(n-f)}^{2}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$, by expending and re-arranging the terms:
((*)( ∗ ) $=a\times[b^{2}(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}+c^{2}(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}+2bc(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}]$
@@
$\lambda_{\max}(\frac{1}{n-f}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal})=\frac{1}{{(n-f)}^{2}}\bigg{(}ab\|g_{A}-g_{B}\|_{2}^{2}+bc\|g_{B}-g_{C}\|_{2}^{2}\\ +ac\|g_{A}-g_{C}\|_{2}^{2}\bigg{)}\qed$
D.5.1 Convex case
Theorem D.13 (Added instabitlity in robust GD - convex case).
Consider the setting described in Section 2 under data poisoning attacks. Let $\mathcal{F}_{\mathcal{Z}}$ denote the set of functions that are $C$-Lipschitz, $L$-smooth and convex on $\mathcal{Z}$. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{2}{L}$. Then there exists a loss function $\ell\in\mathcal{F}_{\mathcal{Z}}$ and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@ D.5.1 Convex case
    Finally, we leverage this construction to derive a lower bound on uniform stability, which matches the upper bound established in Theorem D.10.
@@
$\text{For }\theta\in\mathbb{R},\text{ }z\in[-C,C],\quad\ell(\theta;z)=z\theta,\quad\ell^{\prime}(\theta;z)=z,\quad\ell^{\prime\prime}(\theta;z)=0.$ (17)
@@
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(1,j)}=0,\quad z^{(1,1)}=-C,\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=\frac{1+\psi}{2}C.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=-C.\end{split}$ (18)
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size $n-f$, and for all $t\in\{0,\ldots,T-1\}$,
@@
    which is ensured by construction. In fact, we again recall that the gradients are independent of the parameters and remain constant throughout the optimization process. Next, we demonstrate that any subset composed of a mixture of gradients from ${1,N}$, $F$, and $E$ necessarily yields a larger maximum eigenvalue. This is due to the fact that the vectors $g^{(F)\prime}_{0}$ and $g^{(E)\prime}_{0}$ are collinear but oriented in opposite directions, thereby contributing additively to the overall magnitude. To make this precise, we consider the following two cases:
@@
    $S^{*}_{t}=\{1,N,F\}\Longleftrightarrow\left(n-2f-\frac{2}{m}\right)|g^{(F)}_{t}|^{2}<(n-2f)|g^{(E)}_{t}|^{2}+\frac{2}{m}|g^{(F)}_{t}||g^{(E)}_{t}|.$
    As $|g^{(E)}_{t}|<|g^{(F)}_{t}|$, the above inequality holds if:
@@
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|$ $\geq\ell(\theta_{T};C)=C\theta_{T}=p\gamma C^{2}T.$
D.5.2 Strongly convex case
The structure of the proof with strongly convex objectives closely parallels that of Theorem D.13. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.14 (Added instabitlity in robust GD - strongly convex case).
Consider the setting described in Section 2 under data poisoning attacks. Let $\mathcal{F}_{\mathcal{Z}}$ denote the set of functions that are $C$-Lipschitz, $L$-smooth and $\mu$-strongly convex on $\mathcal{Z}$. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations, $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, $10<c<$ 1, with $\gamma\leq\frac{1}{L}$. Then there exists a loss function $\ell\in\mathcal{F}_{\mathcal{Z}}$ and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@ D.5.2 Strongly convex case
$\ell(\theta;z)=\frac{\mu}{2}{(\theta-z)}^{2},\quad\ell^{\prime}(\theta;z)=\mu(\theta-z),\quad\ell^{\prime\prime}(\theta;z)=\mu$ (19)
@@
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{(x^{(1,1)},y^{(1,1)})\}\cup\{(x^{\prime(1,1)},y^{\prime(1,1)})\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{(x^{(i,1)},y^{(i,1)}),\ldots,(x^{(i,m)},y^{(i,m)})\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(i,j)}=0,\quad z^{(1,1)}=\frac{C}{2\mu},\quad z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=-\frac{1+\psi}{2}\frac{C}{2\mu}.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=\frac{C}{2\mu}.\end{split}$ (20)
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size $n-f$. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
@@
    By Lemma D.12, $S^{*}_{0}=\{1,N,F\}\underset{\mathrm{SMEA}\text{ and }(C1)}{\Longleftarrow}\psi|g^{(F)}_{0}|=\psi\mu\frac{C}{2\mu}<\mu\frac{1+\psi}{2}\frac{C}{2\mu}=|g^{(E)}_{0}|$. We then prove that for all $t\geq 0$, $S^{*}_{t}=\{1,N,F\}$, and the perturbed run converges to the minimizer of the subgroup $F$. For this property to be true, we require:
@@
Provided that there exist $10<c<$ 1 such that $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, this implies: $|\theta_{T}-\theta^{\prime}_{T}|\geq|\theta_{T}|\in\Omega(\frac{pC}{\mu})$. Finally, let define $\chi=-\text{sign}(\theta_{T}+\theta^{\prime}_{T})$, we have:
@@
$\geq\frac{\mu}{2}|\theta_{T}|\frac{C}{\mu}\in\Omega(\frac{pC^{2}}{\mu})$
D.6 Proof of Theorem 3.3 for projected SGD
In this section, we derive a lower bound on the uniform stability of distributed projected SGD with the $\mathrm{SMEA}$ aggregation rule under poisoning attacks. The structure of the proof closely parallels that of Theorem D.13. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.15 (Added instabitlity in robust projected SGD).
Consider the setting described in Section 2 under data poisoning attacks. Let $\mathcal{F}_{\mathcal{Z}}$ denote the set of functions that are $C$-Lipschitz, $L$-smooth and convex on $\mathcal{Z}$. Let $\mathcal{A}=\mathrm{projected-SGD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$. Assume there exist a constant $\tau\geq c>0$ for an arbitrary c, such that $T\geq\tau m$, then there exists a loss function $\ell\in\mathcal{F}_{\mathcal{Z}}$ and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size $n-f$. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
@@
    $S^{*}_{t_{0}-1}\in\underset{}{\text{argmin}}\bigg{\{}\lambda_{\max,t_{0}-1}^{\{1,N,F\}}=\frac{(f+1)(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2},\\ \qquad\qquad\quad\lambda_{\max,t_{0}-1}^{\{1,N,E\}}=\frac{f(n-2f-1)}{{(n-f)}^{2}}\|g^{(E)}_{t_{0}-1}\|_{2}^{2}+\frac{(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2}\\ +\frac{f}{{(n-f)}^{2}}{\left(\|g^{(E)}_{t_{0}-1}\|_{2}+\|g^{(F)}_{t_{0}-1}\|_{2}\right)}^{2}\bigg{\}}.$
@@
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\\ \geq\sum_{t_{0}=1}^{T}\left(1-\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\\ \geq\sum_{t_{0}=1}^{T}\frac{p\gamma L\frac{T-t_{0}+1}{T}}{1+p\gamma L\frac{T-t_{0}+1}{T}}{\left(1-\frac{1}{m}\right)}^{t_{0}-1}$
Where we used that ${\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\leq\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}$ for $T-t_{0}+1\geq 0$ and $-1<-p\gamma Lb^{2}<0$. In fact using $\ln(1+x)\leq x$ for $x>-1$,
$\ln\left[{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}(1+p\gamma Lb^{2}(T-t_{0}+1))\right]$ $=(T-t_{0}+1)\ln(1-p\gamma Lb^{2})$
$\quad+\ln(1+p\gamma Lb^{2}(T-t_{0}+1))$
$\leq-p\gamma Lb^{2}(T-t_{0}+1)+p\gamma Lb^{2}(T-t_{0}+1)$
$=0.$
@@
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\\ \qquad\qquad\geq\frac{p\gamma L}{2}\sum_{t_{0}=1}^{T}(1-\frac{t_{0}-1}{T}){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}(1-{\left(1-\frac{1}{m}\right)}^{T})-\frac{p\gamma L}{2T}\sum_{t_{0}=1}^{T}(t_{0}-1){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}(1-{\left(1-\frac{1}{m}\right)}^{T})-\frac{p\gamma L}{2T}\left[m(m-1)-m^{2}{\left(1-\frac{1}{m}\right)}^{T}(1+\frac{T-1}{m})\right]\hfill\\ =\frac{p\gamma Lm}{2}\left[1-\frac{m-1}{T}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)\right]$
@@
$\mathbb{E}[\|\theta_{T}-\theta^{\prime}_{T}\|_{2}]\geq\|\mathbb{E}[\theta_{T}]\|_{2}\geq\frac{1}{2}(1-\frac{1-e^{-c}}{c})\underbrace{\frac{f+1}{f+\frac{1}{m}}}_{\geq 1}p\gamma CT\in\Omega(p\gamma CT)$
@@
Building on this lower bound for uniform argument stability, we then establish a connection to uniform stability, allowing us to complete the comparison with the upper bound from Theorem D.11.

### v2

@@ C.1 Expectation bounds for honest gradients’ empirical spectral norm
Let $n,f\in\mathbb{N}^{*},f<\frac{n}{2}$ and $g_{A},g_{B},g_{C}$ three collinear vectors in $\mathbb{R}^{d},d\geq 1$. Assume we observe a pool of $n$ vectors, consisting of $\alpha$ copies of $g_{A}$, $\beta$ copies of $g_{B}$ and $\gamma$ copies of $g_{C}$, where $\alpha,\beta,\gamma\in\{0,\ldots,n\}$, $\alpha+\beta+\gamma=n$. We abstract each subset $S\subseteq\{1,\ldots,n\}$, composed of |S|=nf|S|=n-f vectors from our pool of vectors as $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that a+b+c=nfa+b+c=n-f and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. With this notation, we have:
@@
Let $S_{a,b,c}$ with $a\in\{0,\ldots,\alpha\},b\in\{0,\ldots,\beta\},c\in\{0,\ldots,\gamma\}$, such that a+b+c=nfa+b+c=n-f and $S$ is composed by $a$ copies of $g_{A}$, $b$ copies of $g_{B}$ and c copies of $g_{C}$. Then $\overline{g}_{S}=\frac{1}{n-f}(ag_{A}+bg_{B}+cg_{C})$, $g_{A}-\overline{g}_{S}=\frac{b}{n-f}(g_{A}-g_{B})+\frac{c}{n-f}(g_{A}-g_{C})$, $g_{B}-\overline{g}_{S}=\frac{a}{n-f}(g_{B}-g_{A})+\frac{c}{n-f}(g_{B}-g_{C})$ and $g_{C}-\overline{g}_{S}=\frac{a}{n-f}(g_{C}-g_{A})+\frac{b}{n-f}(g_{C}-g_{B})$. We already note that the outer product of vectors $g_{A}$, $g_{B}$ and $g_{C}$ commute because they are collinear. We denote $(*)={(n-f)}^{2}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$, by expending and re-arranging the terms:
$(*)$ $=a\times[b^{2}(g_{A}-g_{B}){(g_{A}-g_{B})}^{\intercal}+c^{2}(g_{A}-g_{C}){(g_{A}-g_{C})}^{\intercal}+2bc(g_{A}-g_{B}){(g_{A}-g_{C})}^{\intercal}]$
@@
$\lambda_{\max}(\frac{1}{n-f}\sum_{i\in S_{a,b,c}}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal})=\frac{1}{{(n-f)}^{2}}\bigg(ab\|g_{A}-g_{B}\|_{2}^{2}+bc\|g_{B}-g_{C}\|_{2}^{2}+ac\|g_{A}-g_{C}\|_{2}^{2}\bigg)$
D.2.1 Convex case
Theorem D.2.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{2}{L}$. Then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@ D.2.1 Convex case
    Finally, we leverage this construction to derive a lower bound on uniform stability, which matches the upper bound established in Theorem 3.3.
@@
$\text{For }\theta\in\mathbb{R},\text{ }z\in[-C,C],\quad\ell(\theta;z)=z\theta,\quad\ell^{\prime}(\theta;z)=z,\quad\ell^{\prime\prime}(\theta;z)=0.$
@@
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(1,j)}=0,\quad z^{(1,1)}=-C,\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=\frac{1+\psi}{2}C.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=-C.\end{split}$
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f, and for all $t\in\{0,\ldots,T-1\}$,
@@
    which is ensured by construction. In fact, we again recall that the gradients are independent of the parameters and remain constant throughout the optimization process. Next, we demonstrate that any subset composed of a mixture of gradients from 1,N{1,N}, $F$, and $E$ necessarily yields a larger maximum eigenvalue. This is due to the fact that the vectors $g^{(F)\prime}_{0}$ and $g^{(E)\prime}_{0}$ are collinear but oriented in opposite directions, thereby contributing additively to the overall magnitude. To make this precise, we consider the following two cases:
@@
[…]
@@
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|\geq\ell(\theta_{T};C)=C\theta_{T}=p\gamma C^{2}T.\qed$
D.2.2 Strongly convex case
The proof structure with strongly convex objectives closely parallels that of Theorem D.2. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.3.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{GD}$, with the $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations, $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, $0<c<1$, with $\gamma\leq\frac{1}{L}$. Then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and $\mu$-stongly convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@ D.2.2 Strongly convex case
$\ell(\theta;z)=\frac{\mu}{2}{(\theta-z)}^{2},\quad\ell^{\prime}(\theta;z)=\mu(\theta-z),\quad\ell^{\prime\prime}(\theta;z)=\mu$
@@
$\begin{split}S=\{D_{1},\ldots,D_{n}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{(x^{(1,1)},y^{(1,1)})\}\cup\{(x^{\prime(1,1)},y^{\prime(1,1)})\}\}\\ \text{ with }\forall i\in\{1,\ldots,n\},&\quad D_{i}=\{(x^{(i,1)},y^{(i,1)}),\ldots,(x^{(i,m)},y^{(i,m)})\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i=1\!\!:z^{(i,j)}=0,\quad z^{(1,1)}=\frac{C}{2\mu},\quad z^{\prime(1,1)}=0.\\ &\quad i\in N=\{2,\ldots,n-2f\},|N|=n-2f-1\!\!:z^{(i,j)}=0.\\ &\quad i\in E=\{n-2f+1,\ldots,n-f\},|E|=f\!\!:z^{(i,j)}=-\frac{1+\psi}{2}\frac{C}{2\mu}.\\ &\quad i\in F=\{n-f+1,\ldots,n\},|F|=f\!\!:z^{(i,j)}=\frac{C}{2\mu}.\end{split}$
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
@@
[…]
@@
Provided that there exist $0<c<1$ such that $T\geq\frac{\ln\left(1-c\right)}{\ln\left(1-\gamma\mu\right)}$, this implies: $|\theta_{T}-\theta^{\prime}_{T}|\geq|\theta_{T}|\in\Omega(\frac{pC}{\mu})$. Finally, let define $\chi=-\text{sign}(\theta_{T}+\theta^{\prime}_{T})$, we have:
@@
$\geq\frac{\mu}{2}|\theta_{T}|\frac{C}{\mu}\in\Omega(\frac{pC^{2}}{\mu})\qed$
D.3 Proof of Theorem 3.4 for SGD
In this section, we derive a lower bound on the uniform stability of distributed $\mathrm{projected-GD}$ with $\mathrm{SMEA}$ under data poisoning. The proof structure closely parallels that of Theorem D.2. We refer to it where appropriate to avoid unnecessary repetition, while explicitly highlighting the key differences.
Theorem D.4.
Consider the setting described in Section 2 under data poisoning. Let $\mathcal{A}=\mathrm{projected-SGD}$ with $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$. Assume there exist a constant $\tau\geq c>0$ for an arbitrary c, such that $T\geq\tau m$, then there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@
In what follows, we denote, for a subset $S\subset\{1,\ldots,n\}$ of size nfn-f. Recall that $\mathrm{SMEA}$ performs averaging over the subset of gradients selected according to the following rule, for all $t\in\{0,\ldots,T-1\}$:
@@
    $S^{*}_{t_{0}-1}\in\underset{}{\text{argmin}}\bigg\{\lambda_{\max,t_{0}-1}^{\{1,N,F\}}=\frac{(f+1)(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2},\\ \qquad\qquad\quad\lambda_{\max,t_{0}-1}^{\{1,N,E\}}=\frac{f(n-2f-1)}{{(n-f)}^{2}}\|g^{(E)}_{t_{0}-1}\|_{2}^{2}+\frac{(n-2f-1)}{{(n-f)}^{2}}\|g^{(F)}_{t_{0}-1}\|_{2}^{2}+\frac{f}{{(n-f)}^{2}}{\left(\|g^{(E)}_{t_{0}-1}\|_{2}+\|g^{(F)}_{t_{0}-1}\|_{2}\right)}^{2}\bigg\}.$
@@
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}$ $\geq\sum_{t_{0}=1}^{T}\left(1-\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}$
$\geq\sum_{t_{0}=1}^{T}\frac{p\gamma L\frac{T-t_{0}+1}{T}}{1+p\gamma L\frac{T-t_{0}+1}{T}}{\left(1-\frac{1}{m}\right)}^{t_{0}-1}$
Where we used that ${\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\leq\frac{1}{1+p\gamma Lb^{2}(T-t_{0}+1)}$ for $T-t_{0}+1\geq 0$ and $1<-p\gamma Lb^{2}<0$. In fact using $\ln(1+x)\leq x$ for x>1x>-1,
$\ln\left[{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}(1+p\gamma Lb^{2}(T-t_{0}+1))\right]$ $=(T-t_{0}+1)\ln(1-p\gamma Lb^{2})+\ln(1+p\gamma Lb^{2}(T-t_{0}+1))$
$\leq-p\gamma Lb^{2}(T-t_{0}+1)+p\gamma Lb^{2}(T-t_{0}+1)=0.$
@@
$\sum_{t_{0}=1}^{T}\left(1-{\left(1-p\gamma Lb^{2}\right)}^{T-t_{0}+1}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\\ \qquad\qquad\geq\frac{p\gamma L}{2}\sum_{t_{0}=1}^{T}\left(1-\frac{t_{0}-1}{T}\right){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)-\frac{p\gamma L}{2T}\sum_{t_{0}=1}^{T}(t_{0}-1){\left(1-\frac{1}{m}\right)}^{t_{0}-1}\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)-\frac{p\gamma L}{2T}\left[m(m-1)-m^{2}{\left(1-\frac{1}{m}\right)}^{T}\left(1+\frac{T-1}{m}\right)\right]\hfill\\ \qquad\qquad=\frac{p\gamma Lm}{2}\left[1-\frac{m-1}{T}\left(1-{\left(1-\frac{1}{m}\right)}^{T}\right)\right]\hfill$
@@
$\mathbb{E}[\|\theta_{T}-\theta^{\prime}_{T}\|_{2}]\geq\|\mathbb{E}[\theta_{T}]\|_{2}\geq\frac{1}{2}(1-\frac{1-e^{-c}}{c})\frac{f+1}{f+\frac{1}{m}}p\gamma CT\in\Omega(p\gamma CT)$
@@
Building on this lower bound for uniform argument stability, we then establish a connection to uniform stability, allowing us to complete the comparison with the upper bound from Theorem 3.3.